OpenStudy (marcelie):

could someone clarify too me the odd , even or neither fucntion... i am so confused

Directrix (directrix):

Start with the definitions. Attached. https://www.mathsisfun.com/algebra/functions-odd-even.html

Directrix (directrix):

Do you have a function which you would like to test?

OpenStudy (marcelie):

yes . one sec

OpenStudy (marcelie):

heres the picture so i got lost

OpenStudy (marcelie):

@directrix

OpenStudy (anonymous):

exponents all odd, polynomial is odd

OpenStudy (marcelie):

can you give me an exmaple

OpenStudy (freckles):

odd positive numbers are 1,3,5,7,9,11,13,15,.... even positive or zero numbers are 0,2,4,6,8.10,12,14,....

OpenStudy (freckles):

\[f(x)=x^2+1 \\ \text{ is even } \\\text{ you can think of it like this } f(x)=x^2+1x^0 \]

OpenStudy (freckles):

2 and 0 are even numbers

OpenStudy (anonymous):

are you sure 0 is even?

OpenStudy (freckles):

it is if i define it that way :p

OpenStudy (freckles):

i define an even number to be any number divisible by 2 0 is divisible by 2 so 0 is even

OpenStudy (anonymous):

\[g(x)=x^3+1\] would be NEITHER is what you meant

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[g(x)=1+3x^3-x^5\\ f(-x)=1-3x^3+x^5\] but \[-(g(x)=-1-3x^3+x^5\neq g(-x)\] so neither

OpenStudy (anonymous):

another example \[f(x)=\frac{x}{x^2-4}\] \[f(-x)=\frac{-x}{(-x)^2-4}=\frac{-x}{x^2-4}=-\left(\frac{x}{x^2-4}\right)=-f(x)\] odd

OpenStudy (marcelie):

i got lost o.o

OpenStudy (marcelie):

so what do u mena if the exponent is even or oddd